OpenStudy (anonymous):

P(x)=x4+2x3-15x2+14x+20 Find all roots of P(x), given that 2+i is a root.

OpenStudy (dumbcow):

x = i+2 x-2 = i (x-2)^2 = -1 x^2 -4x+5 = 0 this gives the factor that divides into P evenly --> (x^2-4x+5)(x^2+6x+4)